# Rapid Gyroscope Calibration: A Deep Learning Approach

### v2

@@ Abstract
Low-cost gyroscope calibration is essential for ensuring the accuracy and reliability of gyroscope measurements. Stationary calibration estimates the deterministic parts of measurement errors. To this end, a common practice is to average the gyroscope readings during a predefined period and estimate the gyroscope bias. Calibration duration plays a crucial role in performance, therefore, longer periods are preferred. However, some applications require quick startup times and calibration is therefore allowed only for a short time. In this work, we focus on reducing low-cost gyroscope calibration time using deep learning methods. We propose a deep-learning framework and explore the possibilities of using multiple real and virtual gyroscopes to improve the calibration performance of single gyroscopes. To train and validate our approach, we recorded a dataset consisting of 169 hours of gyroscope readings, using 24 gyroscopes of two different brands. We also created a virtual dataset consisting of simulated gyroscope readings. The two datasets were used to evaluate our proposed approach. One of our key achievements in this work is reducing gyroscope calibration time by up to 89% using three low-cost gyroscopes.
@@ I Introduction
Inertial navigation systems (INS) are commonly used in manned and autonomous platforms operating in different environments. The INS provides the navigation solution based on its inertial sensor readings [1, 2]. The sensors are arranged in an inertial measurement unit (IMU) consisting of three orthogonal gyroscopes and accelerometers. INSs are popular because of their small size, low cost, high cost-effectiveness, and low power consumption. Yet, the accuracy of the INS depends heavily on the performance of its inertial sensors and their error regime, in particular in the absence of external physical sensor data or information updates[3, 4]. The use of low-cost IMUs further exacerbates this issue because of their large error terms and noise characteristics [5].
To reduce the effect of the sensor error terms on the navigation solution, stationary bias calibration of the inertial sensors is performed before the mission begins. This calibration aims to estimate the deterministic parts of the sensor errors. Once estimated, the errors are removed from the sensor readings during navigation. Calibration approaches can generally be divided into lab and field types. The former requires expensive equipment but it is considered to be more accurate. Lab calibration is essential for ensuring accuracy, consistency, and reliability in sensor measurements. This process requires specialized equipment, including precision turntables and temperature chambers, to meet the stringent demands of various applications. As outlined in [6], the six-position gyroscope calibration method requires a turntable capable of precise angular velocity measurement. Another approach to IMU calibration is temperature compensation, which requires a temperature chamber. Sensor readings can drift because of temperature variations, therefore a temperature sensor is used to record these changes, allowing the creation of a compensation model that adjusts the sensor readings accordingly [7]. A hybrid calibration technique for fiber optic gyroscopes using low-cost IMUs is presented in [8]. In [9], gyroscope calibration is carried out using a pendulum. Additional lab calibration methods include self-calibration through factorization [10], gyroscope denoising [11], and deep learning approaches [12].
@@
In the present research, our goal was to improve stationary gyroscope calibration performance by leveraging DL approaches. To this end, we constructed a simple yet efficient neural network capable of regressing the gyroscope bias. As a single IMU consists of three gyroscopes, we focused on three gyroscope calibration methods. To further enhance our DL calibration approach, we used (a) virtual gyroscope data and (b) multiple gyroscopes (MG) with and without virtual data. We focused only on gyroscopes because of the need for attitude information in zero-order accelerometer calibration. In addition to improvements in calibration accuracy, we sought rapid calibration capability. This approach is particularly beneficial in time-sensitive applications, such as search and rescue operations [37], aerial vehicles [38], and robotics [39], where faster adjustments are critical. Although it may not directly improve precision, it enables quicker deployment and adaptation of the navigation systems, allowing these applications to integrate and use new gyroscopes more rapidly.
To evaluate our proposed approach, we recorded 58.65 hours from 24 gyroscopes of two different brands. In addition, we created a virtual dataset of 110.4 hours. Our results show improvements over the model-based baseline approach both in accuracy and in rapid calibration.
The rest of this paper is organized as follows: Section II describes the model-based calibration approach; section III presents our proposed approaches; section IV provides an analysis and discussion, and section V concludes.
@@ III-B Increasing Input Channels
The basic NN approach uses single gyroscope readings as input to output its bias. We increased the number of input channels, each channel representing a single gyroscope so that the output size corresponds to the number of channels (gyroscopes). As an IMU consists of three orthogonal gyroscopes, with the addition of each IMU, the input channels are increased by three. We sought to train the network with multiple channels and rely on the interconnections between the MG readings to improve the calibration performance. For example, when training the network with data from three gyroscopes, the network input consists of three channels corresponding to the three axes and the network output is the three deterministic bias values. Next, we examined the effect of increasing the number of input channels by incorporating data from additional gyroscopes, resulting in an input of 3N channels, where N is the number of IMUs. Note that the training data increases with the use of addition gyroscopes. If M is the number of samples recorded by each gyro, the training data with consist of $3\ cdotN\ cdotM$ samples.
@@ III-D Neural Network Architecture
It consists of a convolution layer followed by a LeakyReLU activation function [40] and a max-pooling layer. Next, two fully connected layers, with a LeakyReLU activation function between them, process the output features.
@@ IV-A Dataset
We employed two types of gyroscopes for our experiments: (a) Movella Dot [42] and (b) SparkFun [43]. The specifications for these gyroscopes, as provided by the manufacturers, are shown in Table I.
@@
Sensor name Sample rate [Hz] Noise density [mdps/$\sqrt{\text{Hz}}$] Bias stability [$\deg$/h] Movella DOT 120 7 10 SparkFun 130-145 3.8 N/A
We used four Movella DOTs and four SparkFun IMUs in our experiments. During data acquisition, the eight IMUs were placed on a table. The DOT IMUs were operated remotely through a built-in function, allowing for simultaneous recording of all 12 gyroscopes. For the SparkFun IMUs, custom placeholders were designed and 3D printed to accommodate the four IMUs and associated electronics, enabling simultaneous recording and storage of the 12 gyroscope readings. The experimental setup is illustrated in Figure 7.
To achieve an accurate bias estimate from a single recording, a substantial number of samples is necessary to average out the Gaussian mean-zero noise. Consequently, 13,000 samples were recorded for each of the 24 gyroscopes. The bias distribution was estimated by repeating the experiment 100 times. After each experiment, the sensors were powered off for 10 seconds to cool down and return to room temperature.
The SparkFun measurements were conducted at a sample rate between 130 Hz and 145 Hz. For each recording, 13,000 samples were collected over 87 seconds. All SparkFun USB cables were connected to a USB hub, whereas the DOTs were controlled through Bluetooth using a smartphone to start and stop the recordings. For the DOTs, 13,000 samples were collected at 120 Hz over 120 seconds. The total recording time for the DOT gyroscopes was 34 hours and for the SparkFun gyroscopes 24.65 hours. Therefore, our dataset contains a total of 58.65 hours of recordings from 24 gyroscopes.
@@
In addition to the real-world dataset, we generated virtual gyroscope measurements using simulation. The goal of this dataset was twofold: (a) increase the training set data for better results and (b) examine whether less data recorded by real gyroscopes, combined with virtual data, may be used to achieve similar performance in the bias regression task. To create the virtual dataset, we first calculated the bias ground truth values of the 24 gyroscopes (both brands) using (2). The minimum and maximum bias values were identified from the 24 values, for both IMU brands. Next, a bias was derived from a uniform distribution, with the minimum and maximum bias values as limits of the distribution. From the selected bias, we created a normal distribution with a mean equal to the bias and a standard deviation of 0.01 [$\frac{deg}{sec}$]. Finally, we generated gyroscope measurements from this normal distribution. To this end, given a bias value, we generated a sequence of 13,000 virtual measurements for a single recording. In total, for the same bias value, we made 100 recordings. We randomly selected 24 bias values, each representing a virtual gyroscope, and generated 100 recordings for each of them.
@@
Table II shows the duration of all four datasets. Overall, the four datasets contain 169.05 hours of gyroscope measurements from 24 real gyroscopes and 48 virtual ones.
@@
Dataset Train [hours] Test [hours] Total [hours] 1 23.2 1.45 24.65 2 32 2 34 3 46.4 0 46.4 4 64 0 64
@@ IV-B Evaluation Metric and Approach
  • Real2Real: Only real recorded data were used for training, either from Dataset-1 or Dataset-2. In this approach we examined a minimum of three gyroscopes (a single IMU) and a maximum of 12 gyroscopes (4 IMUs).
@@ IV-C Increasing the Number of Input Channels
In this section, we explore the effect of increasing the number of input channels and training data during neural network training on gyroscope calibration accuracy, using the Real2Real approach. Specifically, we analyze the performance of the neural network when using data from Dataset-1. We computed the running RMSE across the test dataset using the model-based approach. Subsequently, we trained the network four times with varying numbers of gyroscopes in the input to the network and training set. We repeated this process twice, once with a 10-second calibration period and again with a 30-second period. Finally, we determined the time difference between the RMSE achieved by the neural network after 10 or 30 seconds and the time required for the model-based approach to reach the same RMSE.
@@ IV-D Increasing the Size of the Training Dataset
We followed the same procedure as in the previous section and examined the influence of the size of the training dataset on the calibration performance. We began the evaluation with three gyroscopes in the training set, and each time, we increased the number of gyroscope recordings by three until we reached 12 gyroscopes. Figure 9 presents the results of our analysis. The blue line in the figure represents the running RMSE average obtained from the model-based approach using the test dataset (the same as in Figure 8). The colored dots show the RMSE in our approach as a function of the number of gyroscopes. We note that as additional gyroscopes are added to the train dataset, the performance improves both for the 10- and the 30-second calibration time. Regardless of the number of gyroscopes, we obtain rapid convergence relative to the model-based approach. For example, when using three gyroscopes, it takes the model-based approach 39 seconds to reach the performance that it takes 10 seconds to achieve in our approach.
Figure 9: Method 2. Increasing the SparkFun training data. The results demonstrate the effect of expanding the training dataset. Our approach shows rapid calibration and improved bias accuracy. All the results were obtained on the same testing dataset.
Figure 10 provides a closer look at the results achieved with the 10-second calibration time. In this case, not only a raid calibration was achieved, but also, for the same calibration time, our calibration network improved upon the accuracy of the model-based approach. For example, when using three gyroscopes, the improvement was 57% and when using 12 gyroscopes, the improvement increased to 88%.
@@ IV-F Approach Robustness
Figure 14 illustrates the outcomes of this analysis. The results attest to the advantages of using MG across different types of gyroscopes. For example, when using 12 DOTs gyroscopes, the model-based approach required 67 seconds to achieve the performance that we obtained using our approach in 10 seconds. With the addition of 24 virtual gyroscopes, an 88% improvement in calibration time was achieved. These results demonstrate that our approach is effective with different types of gyroscopes.
@@ IV-G Summary
As demonstrated, our results consistently indicate significant improvements in calibration time and accuracy. Table III summarizes the results using three gyroscopes (single IMU) and virtual data. In all the approaches, we improved the accuracy of the model-based approach for short calibration times. Additionally, the model-based approach requires more time to achieve the performance that we achieve with a 10-second calibration interval. We demonstrated this performance using two types of IMUs. The addition of virtual data improved both accuracy and calibration time compared to using only three real gyroscopes.

### v3

@@ Abstract
Low-cost gyroscope calibration is essential for ensuring the accuracy and reliability of gyroscope measurements. Stationary calibration estimates the deterministic parts of measurement errors. To this end, a common practice is to average the gyroscope readings during a predefined period and estimate the gyroscope bias. Calibration duration plays a crucial role in performance, therefore, longer periods are preferred. However, some applications require quick startup times and calibration is therefore allowed only for a short time. In this work, we focus on reducing low-cost gyroscope calibration time using deep learning methods. We propose an end-to-end convolutional neural network for the application of gyroscope calibration. We explore the possibilities of using multiple real and virtual gyroscopes to improve the calibration performance of single gyroscopes. To train and validate our approach, we recorded a dataset consisting of 186.6 hours of gyroscope readings, using 36 gyroscopes of four different brands. We also created a virtual dataset consisting of simulated gyroscope readings. The six datasets were used to evaluate our proposed approach. One of our key achievements in this work is reducing gyroscope calibration time by up to 89% using three low-cost gyroscopes. Our dataset is publicly available to allow reproducibility of our work and to increase research in the field.
@@ I Introduction
Inertial navigation systems (INS) are commonly used in manned and autonomous platforms operating in different environments. The INS provides the navigation solution based on its inertial sensor readings [1, 2]. The sensors are arranged in an inertial measurement unit (IMU) consisting of three orthogonal gyroscopes and accelerometers. INSs are popular because of their small size, low cost, high cost-effectiveness, and low power consumption. Yet, the accuracy of the INS depends heavily on the performance of its inertial sensors and their error regime, in particular in the absence of external physical sensor data or information updates[3, 4]. The use of low-cost IMUs further exacerbates this issue because of their large error terms and noise characteristics [5]. Recent research has examined the influence of inertial sensor errors on measurement accuracy and proposed compensation techniques to mitigate their effects [6].
To reduce the effect of the sensor error terms on the navigation solution, stationary bias calibration of the inertial sensors is performed before the mission begins. This calibration aims to estimate the deterministic parts of the sensor errors. Once estimated, the errors are removed from the sensor readings during navigation. Calibration approaches can generally be divided into lab and field types. The former requires expensive equipment but it is considered to be more accurate. Lab calibration is essential for ensuring accuracy, consistency, and reliability in sensor measurements. This process requires specialized equipment, including precision turntables and temperature chambers, to meet the stringent demands of various applications. As outlined in [7], the six-position gyroscope calibration method requires a turntable capable of precise angular velocity measurement. Another approach to IMU calibration is temperature compensation, which requires a temperature chamber. Sensor readings can drift because of temperature variations, therefore a temperature sensor is used to record these changes, allowing the creation of a compensation model that adjusts the sensor readings accordingly [8]. A hybrid calibration technique for fiber optic gyroscopes using low-cost IMUs is presented in [9]. In [10], gyroscope calibration is carried out using a pendulum. Additional in-lab calibration methods include self-calibration through factorization [11], gyroscope denoising [12], IMU hand calibration [13], self-calibration using machine learning [14], and deep learning approaches [15].
@@
This study introduces a neural network-based approach to improve the zero-order calibration of low-cost gyroscopes by leveraging real and virtual data. Our deep learning method significantly reduced calibration time while improving accuracy compared to the baseline model-based approach. Specifically, using a single IMU, our approach demonstrated notable improvements in both accuracy and calibration time. Incorporating virtual data further enhanced these gains, while training with multiple gyroscopes provided additional benefits, particularly in reducing calibration time. These results remained consistent across different gyroscope brands, reinforcing the robustness of our approach. In this work, we integrate our approach with multi-gyroscope (MG) data to significantly reduce gyroscope calibration time while maintaining high accuracy. Unlike previous methods that rely solely on a single gyroscope or extensive calibration durations, our approach leverages multiple real and virtual gyroscopes to enhance bias estimation. Beyond its quantitative improvements, this method has significant practical implications. In applications where rapid calibration is crucial, such as robotics, autonomous vehicles, and search and rescue operations, our approach provides a viable alternative to traditional calibration methods, which require longer stationary periods. The ability to integrate virtual data also reduces the need for large-scale real-world sensor datasets, making this approach more scalable and adaptable to various hardware configurations. By bridging the gap between conventional calibration techniques and data-driven methods, our study contributes to the advancement of rapid and efficient inertial sensor calibration, with broad potential applications in navigation, robotics, and wearable technology. This approach is particularly beneficial in time-sensitive applications, such as search and rescue operations [40], aerial vehicles [41], or air quality measurements using drones [42]. Although it may not directly improve precision, it enables quicker deployment and adaptation of the navigation systems, allowing these applications to integrate and use new gyroscopes more rapidly. To evaluate our proposed approach, we recorded 76.2 hours from 36 gyroscopes of four different brands. In addition, we created a virtual dataset of 110.4 hours. Our real recorded dataset is publicly available at the GitHub repository [^1] https://github.com/ansfl/RapidGyroCalibration , to allow reproducibility of our work and to increase research in the field. Our results show improvements over the model-based baseline approach both in accuracy and in rapid calibration.
The rest of this paper is organized as follows: Section II describes the model-based calibration approach; Section III presents our proposed approaches; Section IV provides an analysis and discussion, and Section V concludes this paper.
@@ III Proposed Approach
In simple terms, our goal is to enhance the accuracy of gyroscope signals by leveraging artificial intelligence (AI). Specifically, we use deep learning, a branch in AI, to predict the inherent errors in the gyroscope measurements and compensate for them, thereby improving overall sensor performance. By applying deep learning techniques, our approach can model complex error patterns that standard calibration methods may not fully capture. This enables precise and reliable motion estimation, which is particularly beneficial for applications such as navigation, robotics, and autonomous systems.
@@ III-B Increasing Input Channels
The basic NN approach uses single gyroscope readings as input to output its bias. We increased the number of input channels, each channel representing a single gyroscope so that the output size corresponds to the number of channels (gyroscopes). As an IMU consists of three orthogonal gyroscopes, with the addition of each IMU, the input channels are increased by three. We sought to train the network with multiple channels and rely on the interconnections between the MG readings to improve the calibration performance. For example, when training the network with data from three gyroscopes, the network input consists of three channels corresponding to the three axes and the network output is the three deterministic bias values. Next, we examined the effect of increasing the number of input channels by incorporating data from additional gyroscopes, resulting in an input of 3N channels, where N is the number of IMUs. Note that the training data increases with the use of addition gyroscopes. If M is the number of samples recorded by each gyro, the training data with consist of $3\cdot N\cdot M$ samples.
@@ III-D Neural Network Architecture
It consists of a convolution layer followed by a LeakyReLU activation function [43] and a max-pooling layer. Next, two fully connected layers, with a LeakyReLU activation function between them, process the output features. Convolutional layers were chosen due to their ability to efficiently extract spatial and temporal patterns from gyroscope data while maintaining computational efficiency. Unlike RNNs or transformers, CNNs require fewer parameters and are less prone to overfitting when applied to relatively short, structured time series. Additionally, LeakyReLU was selected over standard ReLU to prevent issues related to dying neurons, ensuring stable gradient flow during training.
@@ IV-A Dataset
We employed four types of gyroscopes for our experiments: (a) Movella Dot [45], (b) SparkFun [46], (c) NG-IMU [47] and (d) Memsense MS-IMU3025 [48]. The specifications for these gyroscopes, as provided by the manufacturers, are shown in Table I.
@@
Sensor name Sample rate [Hz] Noise density [mdps/$\sqrt{\text{Hz}}$] Bias stability [$\deg$/h] Movella DOT 120 7 10 SparkFun 130-145 3.8 N/A NG 200 N/A N/A Memsense 250 14.8 2.6
We used four Movella DOTs and four SparkFun IMUs in our experiments. After evaluating our approach on those devices, we further tested our approach on the NG and Memsense IMUs to evaluate its generalization. During data acquisition, all 12 IMUs from each brand were placed on a stable table to minimize external disturbances. For synchronization, different methods were applied based on sensor type. Movella DOT sensors were synchronized using built-in Bluetooth functionality, allowing simultaneous activation and recording via a mobile application. For SparkFun, NG, and Memsense IMUs, which do not have built-in synchronization, we designed and 3D-printed custom placeholders to keep the sensors in a fixed position and connected them to a single power hub. A centralized data acquisition script, controlled by Arduino, was implemented to trigger data logging simultaneously across all connected sensors, ensuring time alignment of the recorded gyroscope readings. To account for potential variations in gyroscope performance due to manufacturing inconsistencies, we conducted experiments using multiple units of each IMU type. Despite being from the same manufacturer, individual sensors exhibited slight differences in bias and noise characteristics. By including multiple sensors in our dataset and evaluating performance across different models, we ensured that our approach remains robust to these variations.
@@
To achieve an accurate bias estimate from a single recording, a substantial number of samples is necessary to average out the Gaussian mean-zero noise. Consequently, 13,000 samples were recorded for each of the gyroscopes. The bias distribution was estimated by repeating the experiment 100 times. The training dataset should include a variety of bias values to allow generalization. Consequently, we turned off the device and waited 10 seconds before turning it back on. By doing so, we obtained different bias values in each experiment, which contributed to the training process and allowed for generalization. In addition to the real-world dataset, we generated virtual gyroscope measurements using simulation. The goal of this dataset was twofold: (a) increase the training set data for better results and (b) examine whether less data recorded by real gyroscopes, combined with virtual data, may be used to achieve similar performance in the bias regression task. This process was done only for the DOT and SparkFan. To create the virtual dataset, we first calculated the bias ground truth values of the 24 gyroscopes (both brands) using (2). To ensure that the virtual gyroscope simulation accurately represents real gyroscope characteristics, we derived bias values from the measured uniform distribution of real gyroscopes. The results of this analysis provided an estimate for a reasonable standard deviation range. Each IMU type was assigned a standard deviation value based on the noise characteristics observed in its real recordings, with $0.04$ [$\frac{deg}{sec}$] for the SparkFun gyroscopes. Finally, we generated gyroscope measurements from this normal distribution. To this end, given a bias value, we generated a sequence of 13,000 virtual measurements for a single recording. In total, for the same bias value, we made 100 recordings. We randomly selected 24 bias values, each representing a virtual gyroscope, and generated 100 recordings for each of them.
@@
  • Dataset-5: The dataset contains 100 recordings of each of the 6 NG gyroscopes. Recordings consist of 13,000 measurements lasting 65 seconds. Training set recordings include 8.67 hours, while testing set recordings include 1.08 hours not included in training set recordings.
  • Dataset-6: The dataset contains 100 recordings of each of the 6 Memsense gyroscopes. Recordings consist of 13,000 measurements lasting 52 seconds. Training set recordings include 6.83 hours, while testing set recordings include 0.87 hours not included in training set recordings.
Table II shows the duration of all six datasets. Overall, the six datasets contain 186.6 hours of gyroscope measurements from 36 real gyroscopes and 48 virtual ones.
@@
Dataset Train [hours] Test [hours] Total [hours] 1 23.2 1.45 24.65 2 32 2 34 3 46.4 0 46.4 4 64 0 64 5 8.67 1.08 9.75 6 6.93 0.87 7.8
@@ IV-B Evaluation Metric and Approach
  • Real2Real: Only real recorded data were used for training, either from Dataset-1, Dataset-2, Dataset-5 or Dataset-6. In this approach we examined a minimum of three gyroscopes (a single IMU) and a maximum of 12 gyroscopes (4 IMUs).
@@ IV-C Increasing the Number of Input Channels
In this section, we explore the effect of increasing the number of input channels and training data during neural network training on gyroscope calibration accuracy, using the Real2Real approach. Specifically, we analyze the performance of the neural network when using data from Dataset-1. We computed the running RMSE across the test dataset using the model-based approach. Subsequently, we trained the network four times with varying numbers of gyroscopes in the input to the network and training set. We repeated this process twice, once with a 10-second calibration period and again with a 30-second period. We chose calibration times that represent approximately 10% and 33% of the time required by the model-based approach to converge. From a data-driven point of view, we wanted to give the network significant time to learn and therefore used 10 seconds as a minimum. To meet the inertial requirements, we needed rapid calibration and, thus, limited the maximum time to 30 seconds. Finally, we determined the time difference between the RMSE achieved by the neural network after 10 or 30 seconds and the time required for the model-based approach to reach the same RMSE.
@@ IV-D Increasing the Size of the Training Dataset
We followed the same procedure as in the previous section and examined the influence of the size of the training dataset on the calibration performance. We began the evaluation with three gyroscopes in the training set, and each time, we increased the number of gyroscope recordings by three until we reached 12 gyroscopes. Figure 9 presents the results of our analysis. The blue line in the figure represents the running RMSE average obtained from the model-based approach using the test dataset (the same as in Figure 8). By contrast, the colored dots depict the RMSE of the neural network for varying numbers of gyroscopes. Figures 9, 11, 12, 13, and 14 follow the same concept as Figures 8 and 9, illustrating similar results but using different datasets or methods, as described in Sections IV-B, IV-C, and IV-D. We note that as additional gyroscopes are added to the train dataset, the performance improves both for the 10- and the 30-second calibration time. Regardless of the number of gyroscopes, we obtain rapid convergence relative to the model-based approach. For example, when using three gyroscopes, it takes the model-based approach 39 seconds to reach the performance that it takes 10 seconds to achieve in our approach.
Figure 9: Method 2: Increasing the SparkFun training data. The results demonstrate the effect of expanding the training dataset. Our approach shows rapid calibration and improved bias accuracy. All the results were obtained on the same testing dataset.
Figure 10 provides a closer look at the results achieved with the 10-second calibration time. The blue bars represent the RMSE achieved with our approach using varying numbers of gyroscopes, while the red line corresponds to the model-based approach. This figure provides a closer view of the ten-second calibration results presented in Figure 9. In this case, not only a raid calibration was achieved, but also, for the same calibration time, our calibration network improved upon the accuracy of the model-based approach. For example, when using three gyroscopes, the improvement was 57% and when using 12 gyroscopes, the improvement increased to 88%.
@@ IV-F Approach Robustness
Figure 14 illustrates the outcomes of this analysis. The results attest to the advantages of using MG across different types of gyroscopes. For example, when using 12 DOTs gyroscopes, the model-based approach required 67 seconds to achieve the performance that we obtained using our approach in 10 seconds. With the addition of 24 virtual gyroscopes, an 88% improvement in calibration time was achieved.
@@ IV-G Summary
In [37] a DL approach was used to estimate single gyroscope zero order calibration. While presenting novelty in using DL for gyro calibration, this approach couldn’t achieve better performance for the same calibration time. As demonstrated, using our Real2Real and (Real+ Virtual)2Real approaches, the results consistently indicate significant improvements in calibration time and accuracy. Table III summarizes the results using three gyroscopes (single IMU) and virtual data. In all the approaches, we improved the accuracy of the model-based approach for short calibration times. Additionally, the model-based approach requires more time to achieve the performance that we achieve with a 10-second calibration interval. We demonstrated this performance using two types of IMUs. The addition of virtual data improved both accuracy and calibration time compared to using only three real gyroscopes.
@@
TABLE V: Summary of calibration results using three real gyroscopes without virtual data in a 10-second calibration time.
IMU type Number of gyroscopes Three gyroscopes (single IMU) Improvement using our approach Calibration time (same performance) [%] Accuracy improvement (same calibration time) [%] SparkFan 9 Real 82 87 DOT 9 Real 89 82 NG 6 Real 83 98 Memsense 6 Real 75 75
Table V presents the results for the four IMU types using the Real2Real Method with increasing training data. The results demonstrate our approach’s versatility across different IMU types. Three of the IMU manufacturers examined with our approach achieved calibration time and accuracy results of 80% or higher. Using the Memsense IMU the improvement was 75%. These results highlight our method’s effectiveness.